Given the input current 15A and the output voltage of 100V for a step-up transformer
having 90% efficiency, find the output power and the voltage in the secondary if the
output current is 3 A.
1
Expert's answer
2020-02-06T08:29:13-0500

a) By the definition of the efficiency we have:


"\eta = \dfrac{P_{output}}{P_{input}} \cdot 100 \%,"

here, "P_{input}" is the input power in the primary (input) winding of the transformer, "P_{output}" is the output power in the secondary (output) winding of the transformer, "\eta = 90 \%" is the efficiency of the transformer.

We can find the input power from the formula:


"P_{input} = I_{input}V_{input},"

here, "I_{input} = 15A" is the input current in the primary winding of the transformer and "V_{input} = 100V" is the input voltage in the primary winding of the transformer.

Substituting the expression for "P_{input}" into the formula for the efficiency we can calculate the output power in the secondary winding of the transformer:


"P_{output} = \dfrac{\eta}{100 \%} I_{input}V_{input} = \dfrac{90 \%}{100 \%} \cdot 15A \cdot 100V = 1350W."

b) We can find the output voltage in the secondary winding of the transformer from the formula:


"P_{output} = I_{output}V_{output},""V_{output} = \dfrac{P_{output}}{I_{output}} = \dfrac{1350W}{3A} = 450V."

Answer:

a) "P_{output} = 1350W."

b) "V_{output} = 450V."